OpenStudy (anonymous):

Find the slope of the line when the points are (6,6) and (7,0)

OpenStudy (bahrom7893):

Slope = (y-y1)/(x-x1)

OpenStudy (bahrom7893):

y = 6, y1 = 0 x = 6, x1 = 7

OpenStudy (saifoo.khan):

\[m = \frac {0-6}{7-6} \]

OpenStudy (saifoo.khan):

therefore: -6 --- 1 final slope = -6.

OpenStudy (bahrom7893):

The answer is -6

OpenStudy (bahrom7893):

It's supposed to be: M = (6-0)/(6-7) = 6/-1=-6

OpenStudy (anonymous):

what would the answer be if you flipped the points around?

OpenStudy (bahrom7893):

it would be the same

OpenStudy (saifoo.khan):

Yes, it will

OpenStudy (bahrom7893):

doesn't matter which point u pick to use for y1 or y

OpenStudy (bahrom7893):

just make sure to flip the x1 and x as well

OpenStudy (saifoo.khan):

i mean, yes it will remain the same

OpenStudy (saifoo.khan):

Megan Fox is here! :O

OpenStudy (bahrom7893):

WOO!! Is that why you put that pic up?

OpenStudy (saifoo.khan):

Haha, No.